1) Larkspur, Inc. issues 8,700 shares of $101 par value preferred stock for cash at $118 per share.

Journalize the issuance of the preferred stock.

2)The stockholders’ equity section of Culver Corporation’s balance sheet consists of common stock ($8 par) $1,040,000 and retained earnings $500,000. A 10% stock dividend (13,000 shares) is declared when the market price per share is $18.

(a) Show the before-and-after effects of the dividend on the components of stockholders’ equity.
Answer #1
1
Debit Credit
Cash 1026600 =8700*118
    Preferred Stock 878700 =8700*101
    Paid in Capital in excess of par-Preferred Stock 147900
To record the issuance of the preferred stock.
2
Before dividend After dividend
Stockholders’ equity
Paid-in capital
Common stock 1040000 1144000
Paid in Capital In excess of par-Common Stock 0 130000
Total paid-in capital 1040000 1274000
Retained earnings 500000 266000
Total stockholders’equity 1540000 1540000
Workings:
Value of Stock Dividend 234000 =13000*18
Common stock increase 104000 =13000*8
Paid in Capital In excess of par-Common Stock 130000 =13000*(18-8)
Retained earnings 266000 =500000-234000